OpenStudy (anonymous):

How would I simplify this into a fraction? ( 2x+6)=( 8+10x)

OpenStudy (anonymous):

what does it mean "simplify in to a fraction"?

OpenStudy (anonymous):

here's the real question and that was all i got so far What is the solution of the equation? Eliminate any extraneous solutions. (-2x+6) 1/5 = (-8+10x) 1/5 A. 7/6 B.-1/4 C. 2/3 D. 6/7

OpenStudy (anonymous):

\[ 2x+6= 8+10x\] oooh i see, it means "solve for \(x\)" got it

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

but maybe you dropped a couple minus signs should it be \[-2x+6= -8+10x\]??

OpenStudy (anonymous):

anyway you can post the original question? a picture or something?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

this is the original question What is the solution of the equation? Eliminate any extraneous solutions. (-2x+6) 1/5 = (-8+10x) 1/5 A. 7/6 B.-1/4 C. 2/3 D. 6/7

OpenStudy (anonymous):

THANKS DIALA

OpenStudy (anonymous):

\[\frac{-2x+6}{5} = \frac{-8+10x}{5} \]?